OpenStudy (ttp):

A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t2 + 32t + 90 The average rate of change of f(t) from t = 4 seconds to t = 6 seconds is _____ feet per second.

OpenStudy (anonymous):

Solve 2x2−12x+5=0 pattern match this equation with ax2+bx+c=0 we see a=2, b= -12 , and c= 5 use those numbers to replace a,b, and c in the quadratic formula x=−b±b2−4ac−−−−−−−√2a notice we use the discriminant, which you know how to find. b2−4ac becomes 144 - 4*2*5= 144-40= 104 x=12±104−−−√2⋅2124±226−−√4x=3±1226−−√ if we put one of these solutions into the original equation we will get zero (just as we should)

OpenStudy (ttp):

?